# MantisScore: Building Automatic Metrics to Simulate Fine-grained Human Feedback for Video Generation

### v1

@@ Abstract
The recent years have witnessed great advances in video generation. However, the development of automatic video metrics is lagging significantly behind. None of the existing metric is able to provide reliable scores over generated videos. The main barrier is the lack of large-scale human-annotated dataset. In this paper, we release VideoFeedback, the first large-scale dataset containing human-provided multi-aspect score over 37.6K synthesized videos from 11 existing video generative models. We train MantisScore (initialized from Mantis) based on VideoFeedback to enable automatic video quality assessment. Experiments show that the Spearman correlation between MantisScore and humans can reach 77.1 on VideoFeedback-test, beating the prior best metrics by about 50 points. Further result on other held-out EvalCrafter, GenAI-Bench, and VBench show that MantisScore has consistently much higher correlation with human judges than other metrics. Due to these results, we believe MantisScore can serve as a great proxy for human raters to (1) rate different video models to track progress (2) simulate fine-grained human feedback in Reinforcement Learning with Human Feedback (RLHF) to improve current video generation models.
@@
MantisScore: Building Automatic Metrics to Simulate Fine-grained Human Feedback for Video Generation
@@
https://tiger-ai-lab.github.io/MantisScore/
Figure 1: Construction process of VideoFeedback dataset and illustration of MantisScore.
@@ 1 Introduction
Since obtaining large-scale human feedback is highly costly, we can try to approximate human-provided scores with model-based metrics. To this end, our work can be divided into two parts: (1) curating VideoFeedback, the first large-scale dataset containing human-provided scores for 37.6K videos, (2) training MantisScore on VideoFeedback, which is an automatic video metric to simulate human feedback.
@@
To build the video evaluator, we select Mantis-Idefics2-8B Jiang et al. (2024a) as our main backbone model due to its superior ability to handle multi-image and video content, accommodating up to 128 video frames and supporting native resolution. After fine-tuning Mantis on VideoFeedback-train, we get our video evaluator, MantisScore. Experiments show that we achieve a Spearman correlation of $77.1$ on VideoFeedback-test and $59.5$ on EvalCrafter Liu et al. (2023b) for the text-to-video alignment aspect, surpassing the best baseline by $54.1$ and $4.4$ respectively. The pairwise comparison accuracy gets $78.5$ on GenAI-Bench Jiang et al. (2024b) video preference part, and $72.1$ in average on 5 aspects of VBench Huang et al. (2023), surpassing the previous best baseline by $11.4$ and $9.6$ respectively. Additional ablation studies with different backbone models confirmed that the Mantis-based metric provides a gain of $12.1$ compared to using the Idefics2-based metric. Due to the significant improvement, we believe that MantisScore can serve as the reliable metrics for future video generative models.
@@ 2.3 RLHF in image/video generation
In the recent years, reinforcement learning from human feedback (RLHF) has emerged as a significant approach in enhancing the performance of image/video generative models. Numerous studies have focused on training reward models with large datasets of image-text pairs, such as HPSv2 Wu et al. (2023b), ImageReward Xu et al. (2023), RichHF-18K Liang et al. (2023), or video-text pairs like T2V-Score Wu et al. (2024), while some research efforts have combined existing reward models, such as T2V-Turbo Li et al. (2024b), to simulate human feedback. Utilizing these reward models or feedback simulators, diverse methods have been proposed to align the output of visual generative models with human preferences, including RL-based methods (Fan et al. (2024), Zhang et al. (2024)) and reward fine-tuning methods (Clark et al. (2023), Li et al. (2024a), Yuan et al. (2023)). Additionally, some works adopts data distillation to fine-tune diffusions models on high-quality data, while others like Diffusion-DPO Wallace et al. (2023), extend Direct Preference Optimization (DPO) to train diffusion models based on preference data. Our MantisScore aims to approximate human feedback, which is expected to be beneficial in enhancing video generative models with RLHF methods like PPO or DPO.
@@ 4 Experiments
In this section, we describe our experiment setup, including baseline methods for video evaluation, and evaluation benchmarks for video evaluation. We also discuss the training details of MantisScore, and the analysis of our experiment results.
@@ 4.2 Evaluation Benchmarks
We have included the following benchmarks to evaluate the ability of MantisScore and the above-mentioned baselines on evaluating model generation results.
@@
Table 6: Spearman’s Correlation ($\rho$) of MantisScore on EvalCrafter Liu et al. (2023b).
@@
Table 7: Ablation study on the base model and scoring function for MantisScore. "" means that we take the average of Spearman correlation or pairwise accuracy across the multiple aspects of the benchmark. The highest numbers are bold for each benchmark, and the second are underlined.
@@ 4.3 Training Details
For MantisScore, We use two scoring methods: generative scoring and regression scoring. Generative scoring involves training the model to output fixed text forms, from which aspect scores are extracted using regular expressions. These scores are integers corresponding to human annotation scores. In contrast, regression scoring replaces the language model head with a linear layer that outputs 5 logits representing scores for each aspect. Regression scoring is trained using MSE loss.
@@ 4.4 Results
MantisScore achieves the SoTA performance
On the VideoFeedback-test, MantisScore gets an average of $54.1$ improvements on all the five aspects compared to the baseline GPT-4o. What’s more, on the EvalCrafter benchmark, MantisScore (reg) has $4.4$ improvements on text-to-video alignment. For pairwise preference comparison, MantisScore also gets $78.5$ accuracy on GenAI-Bench, surpassing the second-best Gemini-1.5-Flash by $11.4$ points. on the Vbench, our model archives the highest pairwise accuracy on 4 out of 5 aspects from VBench, with an average of $16.1$ improvements.
@@ 4.5 Ablation Study
We conducted an ablation study on the base model selection and scoring types by training different variants on VideoFeedback. Results shown in Table 7.
@@
The primary difference between regression scoring and generative scoring is that regression scoring can give more fine-grained scores instead of just the four labels. Results on EvalCrafter, GenAI-Bench, and VBench all indicate that using regression scoring can consistently improve the Spearman correlation or the pairwise comparison accuracy. For example, on GenAI-Bench, MantisScore (reg) achieves $78.5$ accuracy, which is higher than the $59.0$ of the MantisScore (gen). The results are similar for the other benchmarks. We thus conclude that regression scoring with more fine-grained scores is a better choice.
@@ 5 Conclusion
In this paper, we introduce MantisScore, which is trained on our meticulously curated dataset VideoFeedback for video evaluation and can serve as good simulator for human feedback on generated videos. We hired 20 expert raters to annotate the 37.6K videos generated from 11 popular text-to-video generative models across 5 key aspects, Visual Quality, Temporal Consistency, Dynamic Degree, Text-to-Video Alignment and Factual Consistency. Our IAA match ratio gets more than 60%. We test the performance of MantisScore using Spearman correlation on VideoFeedback-test and EvalCrafter, and using pairwise comparison accuracy on GenAI-Bench and VBench. The results show that MantisScore consistently gets the best performance, surpassing the powerful baseline GPT-4o and Gemini 1.5 Flash/Pro by a large margin. Our work highlights the importance of using MLLM for video evaluation and demonstrates the future prospects of simulating human scores or feedback in generative tasks, due to its rich world knowledge and the high-quality rating dataset with fine-grained and multiple dimensions.
@@ Appendix B Risks and Limitation
Although we have designed systematic pipelines to recruit expert raters and annotate the video evaluation scores, we still find out that some annotations contain errors and may harm the overall quality of the dataset. Our IAA score computation is only based on a small number of trial examples and, thus might not represent the actual IAA of the whole annotations.
Besides, while MantisScore is proven to be able to effectively give reasonable scores on our defined five aspects, it can still sometimes output wrong scores that do not match our expectations. We admit this drawback and list that as one of our future works.
@@ Appendix G Prompting Template
Table 9: Prompting template in generation format used for MantisScore training and the MLLM prompting baselines
@@
Table 10: Prompting template used for the MLLM prompting baseline and MantisScore training
@@ Appendix H Feature-based Baselines Discretization
Dynamic Degree MSE-$\uparrow$ [0,100] [100,1000) [1000,3000) [3000$\infty$)

### v2

# VideoScore: Building Automatic Metrics to Simulate Fine-grained Human Feedback for Video Generation

@@ Abstract
The recent years have witnessed great advances in video generation. However, the development of automatic video metrics is lagging significantly behind. None of the existing metric is able to provide reliable scores over generated videos. The main barrier is the lack of large-scale human-annotated dataset. In this paper, we release VideoFeedback, the first large-scale dataset containing human-provided multi-aspect score over 37.6K synthesized videos from 11 existing video generative models. We train VideoScore (initialized from Mantis) based on VideoFeedback to enable automatic video quality assessment. Experiments show that the Spearman correlation between VideoScore and humans can reach 77.1 on VideoFeedback-test, beating the prior best metrics by about 50 points. Further result on other held-out EvalCrafter, GenAI-Bench, and VBench show that VideoScore has consistently much higher correlation with human judges than other metrics. Due to these results, we believe VideoScore can serve as a great proxy for human raters to (1) rate different video models to track progress (2) simulate fine-grained human feedback in Reinforcement Learning with Human Feedback (RLHF) to improve current video generation models.
@@
VideoScore: Building Automatic Metrics to Simulate Fine-grained Human Feedback for Video Generation
@@
https://tiger-ai-lab.github.io/VideoScore/
Figure 1: Construction process of VideoFeedback dataset and illustration of VideoScore.
@@ 1 Introduction
Since obtaining large-scale human feedback is highly costly, we can try to approximate human-provided scores with model-based metrics. To this end, our work can be divided into two parts: (1) curating VideoFeedback, the first large-scale dataset containing human-provided scores for 37.6K videos, (2) training VideoScore on VideoFeedback, which is an automatic video metric to simulate human feedback.
@@
To build the video evaluator, we select Mantis-Idefics2-8B Jiang et al. (2024a) as our main backbone model due to its superior ability to handle multi-image and video content, accommodating up to 128 video frames and supporting native resolution. After fine-tuning Mantis on VideoFeedback-train, we get our video evaluator, VideoScore. Experiments show that we achieve a Spearman correlation of $77.1$ on VideoFeedback-test and $59.5$ on EvalCrafter Liu et al. (2023b) for the text-to-video alignment aspect, surpassing the best baseline by $54.1$ and $4.4$ respectively. The pairwise comparison accuracy gets $78.5$ on GenAI-Bench Jiang et al. (2024b) video preference part, and $72.1$ in average on 5 aspects of VBench Huang et al. (2023), surpassing the previous best baseline by $11.4$ and $9.6$ respectively. Additional ablation studies with different backbone models confirmed that the Mantis-based metric provides a gain of $12.1$ compared to using the Idefics2-based metric. Due to the significant improvement, we believe that VideoScore can serve as the reliable metrics for future video generative models.
@@ 2.3 RLHF in image/video generation
In the recent years, reinforcement learning from human feedback (RLHF) has emerged as a significant approach in enhancing the performance of image/video generative models. Numerous studies have focused on training reward models with large datasets of image-text pairs, such as HPSv2 Wu et al. (2023b), ImageReward Xu et al. (2023), RichHF-18K Liang et al. (2023), or video-text pairs like T2V-Score Wu et al. (2024), while some research efforts have combined existing reward models, such as T2V-Turbo Li et al. (2024b), to simulate human feedback. Utilizing these reward models or feedback simulators, diverse methods have been proposed to align the output of visual generative models with human preferences, including RL-based methods (Fan et al. (2024), Zhang et al. (2024)) and reward fine-tuning methods (Clark et al. (2023), Li et al. (2024a), Yuan et al. (2023)). Additionally, some works adopts data distillation to fine-tune diffusions models on high-quality data, while others like Diffusion-DPO Wallace et al. (2023), extend Direct Preference Optimization (DPO) to train diffusion models based on preference data. Our VideoScore aims to approximate human feedback, which is expected to be beneficial in enhancing video generative models with RLHF methods like PPO or DPO.
@@ 4 Experiments
In this section, we describe our experiment setup, including baseline methods for video evaluation, and evaluation benchmarks for video evaluation. We also discuss the training details of VideoScore, and analyze our experiment results.
@@ 4.2 Evaluation Benchmarks
We have included the following benchmarks to evaluate the ability of VideoScore and the above-mentioned baselines on evaluating model generated videos to see their correlation with human raters.
@@
Table 6: Spearman’s Correlation ($\rho$) of VideoScore on EvalCrafter Liu et al. (2023b).
@@
Table 7: Ablation study on the base model and scoring function for VideoScore. "" means that we take the average of Spearman correlation or pairwise accuracy across the multiple aspects of the benchmark. The highest numbers are bold for each benchmark, and the second are underlined.
@@ 4.3 Training Details
For VideoScore, We use two scoring methods: generative scoring and regression scoring. Generative scoring involves training the model to output fixed text forms, from which aspect scores are extracted using regular expressions. These scores are integers corresponding to human annotation scores. In contrast, regression scoring replaces the language model head with a linear layer that outputs 5 logits representing scores for each aspect. Regression scoring is trained using MSE loss.
@@ 4.4 Results
VideoScore achieves the SoTA performance
On the VideoFeedback-test, VideoScore gets an average of $54.1$ improvements on all the five aspects compared to the baseline GPT-4o. What’s more, on the EvalCrafter benchmark, VideoScore (reg) has $4.4$ improvements on text-to-video alignment. For pairwise preference comparison, VideoScore also gets $78.5$ accuracy on GenAI-Bench, surpassing the second-best Gemini-1.5-Flash by $11.4$ points. on the Vbench, our model archives the highest pairwise accuracy on 4 out of 5 aspects from VBench, with an average of $16.1$ improvements.
@@ 4.5 Ablation Study
We conducted an ablation study on the base model selection and scoring types by training different variants on VideoFeedback. Results of the ablation study are shown in Table 7.
@@
The primary difference between regression scoring and generative scoring is that regression scoring can give more fine-grained scores instead of just the four labels. Results on EvalCrafter, GenAI-Bench, and VBench all indicate that using regression scoring can consistently improve the Spearman correlation or the pairwise comparison accuracy. For example, on GenAI-Bench, VideoScore (reg) achieves $78.5$ accuracy, which is higher than the $59.0$ of the VideoScore (gen). The results are similar for the other benchmarks. We thus conclude that regression scoring with more fine-grained scores is a better choice.
@@ 5 Conclusion
In this paper, we introduce VideoScore, which is trained on our meticulously curated dataset VideoFeedback for video evaluation and can serve as good simulator for human feedback on generated videos. We hired 20 expert raters to annotate the 37.6K videos generated from 11 popular text-to-video generative models across 5 key aspects, Visual Quality, Temporal Consistency, Dynamic Degree, Text-to-Video Alignment and Factual Consistency. Our IAA match ratio gets more than 60%. We test the performance of VideoScore using Spearman correlation on VideoFeedback-test and EvalCrafter, and using pairwise comparison accuracy on GenAI-Bench and VBench. The results show that VideoScore consistently gets the best performance, surpassing the powerful baseline GPT-4o and Gemini 1.5 Flash/Pro by a large margin. Our work highlights the importance of using MLLM for video evaluation and demonstrates the future prospects of simulating human scores or feedback in generative tasks, due to its rich world knowledge and the high-quality rating dataset with fine-grained and multiple dimensions.
@@ Appendix B Risks and Limitation
Although we have designed systematic pipelines to recruit expert raters and annotate the video evaluation scores, we still find out that some annotations contain errors and may harm the overall quality of the dataset. Our IAA score computation is only based on a small number of trial examples and, thus might not represent the actual IAA of the whole annotations. Besides, while VideoScore is proven to be able to effectively give reasonable scores on our defined five aspects, it can still sometimes output wrong scores that do not match our expectations. We admit this drawback and list that as one of our future works.
@@ Appendix G Prompting Template
Table 9: Prompting template in generation format used for VideoScore training and the MLLM prompting baselines
@@
Table 10: Prompting template used for the MLLM prompting baseline and VideoScore training
@@ Appendix H Feature-based Baselines Discretization
Dynamic Degree MSE-$\uparrow$ [0,100) [100,1000) [1000,3000) [3000$\infty$)